With a short time remaining in the day, a delivery driver has time to make deliveries at 3 locations among the 8 locations remaining. How many different routes are possible?
Answer #1

Nunber of ways of choosing 3 out of 9 destinations and arranging ( where order of delivery is important ) is,

8P3 = 8! / (8 - 3)! = 8 * 7 * 6 * 5! / 5! = 8 * 7 * 6 = 336

Therefore, there are 336 different routes are possible.